The seeds in bush bean pods are each the product of an independent fertilization event. Green seed color is dominant to white4 green and 2 white seeds? Express your answer to three decimal places (example: 0.123). VO AED A 0 2 ? Submit Previous Answeall green seeds? Express your answer to three decimal places (example: 0.123). O ALQ * R 0 2 ? Submit Request Answerat least 1 white seed? Express your answer to three decimal places (example: 0.123). IVO AQ R o 2 ? Submit Request Answer
Answer #1

Let the allele of green seed color be G

Let the allele of white seed color be g

As the plant is heterozygous green its genotype is Gg.

When this plant is self crossed we get: Gg x Gg

G

g

G

GG

(green seeds)

Gg

(green seeds)

g

Gg

(green seeds)

gg

(white seeds)

So, from the punnett square it is clear that the,

probability of getting green seeds= 3/4

probability of getting white seeds= 1/4.

A)

Probability of 4 green seeds and 2 white seeds

We multiply the probability of green seeds 4 times and white seeds 2 times.

=3/4 x 3/4 x 3/4 x 3/4 x 1/4 x 1/4

= 0.0197 = 0.020

B)

Probability of all green seeds

We multiply the probability of green sees 6 times

= 3/4 x 3/4 x 3/4 x 3/4 x 3/4 x 3/4

= 0.1779 = 0.180

C)

Probability of at least one white seed

= 1 - Probability of all green seeds ( As in all cases other than in cases of all green seeds, there will be at least one white seed)

= 1 - 0.180

= 0.820